# What Can String Probability Tell Us About Grammaticality?

### v1

@@ 6.1 Transformations of probability
The ability of this metric to separate grammatical and ungrammatical sentences has been explicitly investigated (Lau et al., 2017; Leivada et al., 2024a), and is also the implicit standard for minimal pair comparisons (e.g,. Warstadt et al., 2020).
@@
This is equivalent to the “Norm LP (Sub)” metric proposed by Lau et al. (2017), or, equivalently, SLOR without the length-normalization factor.
@@
where ${p}({{w}_{n}})$ is the unigram (i.e., frequency) estimate of word ${w}_{n}$}. This metric is equivalent to the Syntactic Log-Odds Ratio (SLOR) proposed by Pauls and Klein (2012), which has also been investigated by Lau et al. (2017), although the connection to PMI has not been previously established.
@@ 6.2 Evaluation materials
To test ˜3, we no longer need paired sentences (as was the case for Predictions 1 and 2), but instead simply need large sets of grammatical and ungrammatical sentences from which to compute the relevant metrics. We evaluate models on five English datasets that contrast ungrammatical and grammatical sentences: the three minimal-pair datasets used to test ˜1 (BLiMP, SCaMP, and SyntaxGym), the LI dataset used to test ˜2, and CoLA (Warstadt et al., 2019). [^6] Here, we focus on English due to the computational demands of token-frequency estimation for SLOR. See LABEL:tab:datasets for a summary.
@@ 6.3 Computing separation
We evaluate the same two LMs used to evaluate the English datasets for Predictions 1 and 2: GPT-2 and Llama-3-70B (see LABEL:tab:models). To obtain token frequency measurements for SLOR, we sought to estimate the distribution of tokens in each model’s training data. [^7] Our SLOR metric is computed over tokens, not words. Since we do not have access to this data, we used the Huggingface FineWeb dataset (Penedo et al., 2024) as a representative sample of a high-quality Internet text corpus. We used each model to tokenize 1 million items from the sample-10BT sample of FineWeb, and then used this sample to estimate the token frequency distribution for each model.
@@ 7 Discussion
It is uncontroversial that string probability is not the same as grammaticality. But that does not mean that string probability cannot reveal information about a probabilistic model’s underlying grammatical knowledge. Here, we argued that these probabilities are determined by a combination of the probability of the message and the grammaticality of the string. Our theoretical framework shows that some prior critiques of minimal-pair analysis—e.g., that probability does not robustly separate grammatical and ungrammatical strings (Leivada et al., 2024a, b)—fall out of simple assumptions about the generative process underlying linguistic corpus data.
@@
An offshoot of our analysis is that studies of grammaticality in LMs that use minimal pairs that are not tightly controlled (e.g., Vázquez Martínez et al., 2023) risk underestimating the grammatical competence of models by failing to control for the influence of M{M}. In other words, a model could seemingly not differentiate between grammatical and ungrammatical strings, if the messages across grammatical and ungrammatical strings are not well-controlled. As we have argued here, this does not necessarily imply that the model has not learned generalizations about grammatical rules. If we are interested in isolating the model’s sensitivity to grammaticality, we have to use carefully designed procedures to factor out M{M}.
@@
One such procedure is minimal pair string comparisons, where the members of the minimal pair are closely matched in M{M} but are hypothesized to differ in G{G}. While controlled minimal pair comparisons are hardly new in NLP (e.g., Marvin and Linzen, 2018; Futrell et al., 2019), we have provided new theoretical grounding for these practices. In addition, our work lays the foundation for using computational techniques to isolate the effects of M{M} and G{G}, which has been explored in recent work (Stańczak et al., 2024).
@@
Our analyses also raise new questions regarding LMs’ grammatical knowledge. The poor separation achieved by state-of-the-art LMs in Section˜6 feels counterintuitive: if LMs virtually always produce grammatical strings (under standard sampling procedures), then why is there so much overlap between the probabilities assigned to grammatical and ungrammatical strings? This tension between discriminative failures and generative abilities could be seen as a specific realization of the “generative AI paradox” (West et al., 2024), and also connects to recent work demonstrating that language identification is impossible except in highly constrained cases (Gold, 1967; Angluin, 1980), whereas language generation is possible for any countable list of languages (Kleinberg and Mullainathan, 2024).
@@
Leivada et al. (2024b) argue that comparing isolated acceptability judgments in humans against minimal-pair probability differences in models is not comparing “apples with apples”, and thus unfair. We hope that our theoretical model can bring greater clarity to the issue of what makes a fair comparison. When comparing the (cognitive) abilities of two groups—humans and models, younger and older children, or even two different animal species—we maintain that the researcher must design assessments with that group’s (cognitive) computational architecture in mind. Applying the same evaluation method, which might impose auxiliary challenges on one group but not the other, can artificially increase apparent intergroup differences (Firestone, 2020; Lampinen, 2024; Hu and Frank, 2024). For example: the intelligence of a squirrel should not be judged based on its ability to solve a Rubik’s cube. [^8] This example is due to Andrew Lampinen. Similarly, using metalinguistic judgments or isolated string probability as a window into grammaticality ignores the reality of what LMs are, and what they are trained to do—namely, to maximize the probability of strings from a corpus. More broadly, linguistic theory continues to suggest new ways to evaluate LMs, just as modern LMs provide new tools for studying the relationship between probability and grammaticality.

### v2

@@ 6.1 Transformations of probability
The ability of this metric to separate grammatical and ungrammatical sentences has been explicitly investigated (lau_grammaticality_2017; leivada_evaluating_2024), and is also the implicit standard for minimal pair comparisons (e.g,. warstadt_blimp_2020).
@@
This is equivalent to the “Norm LP (Sub)” metric proposed by lau_grammaticality_2017, or, equivalently, SLOR without the length-normalization factor.
@@
where ${p}({{w}_{n}})$ is the unigram (i.e., frequency) estimate of word ${w}_{n}$}. This metric is equivalent to the Syntactic Log-Odds Ratio (SLOR) proposed by pauls_large-scale_2012, which has also been investigated by lau_grammaticality_2017, although the connection to PMI has not been previously established.
@@ 6.2 Evaluation materials
To test ˜3, we no longer need paired sentences (as was the case for Predictions 1 and 2), but instead simply need large sets of grammatical and ungrammatical sentences from which to compute the relevant metrics. We evaluate models on five English datasets that contrast ungrammatical and grammatical sentences: the three minimal-pair datasets used to test ˜1 (BLiMP, SCaMP, and SyntaxGym), the LI dataset used to test ˜2, and CoLA (warstadt_neural_2019). [^6] Here, we focus on English due to the computational demands of token-frequency estimation for SLOR. See LABEL:tab:datasets for a summary.
@@ 6.3 Computing separation
We evaluate the same two LMs used to evaluate the English datasets for Predictions 1 and 2: GPT-2 and Llama-3-70B (see LABEL:tab:models). To obtain token frequency measurements for SLOR, we sought to estimate the distribution of tokens in each model’s training data. [^7] Our SLOR metric is computed over tokens, not words. Since we do not have access to this data, we used the Huggingface FineWeb dataset (penedo_fineweb_2024) as a representative sample of a high-quality Internet text corpus. We used each model to tokenize 1 million items from the sample-10BT sample of FineWeb, and then used this sample to estimate the token frequency distribution for each model.
@@ 7 Discussion
It is uncontroversial that string probability is not the same as grammaticality. But that does not mean that string probability cannot reveal information about a probabilistic model’s underlying grammatical knowledge. Here, we argued that these probabilities are determined by a combination of the probability of the message and the grammaticality of the string. Our theoretical framework shows that some prior critiques of minimal-pair analysis—e.g., that probability does not robustly separate grammatical and ungrammatical strings (leivada_evaluating_2024; leivada_reply_2024)—fall out of simple assumptions about the generative process underlying linguistic corpus data.
@@
An offshoot of our analysis is that studies of grammaticality in LMs that use minimal pairs that are not tightly controlled (e.g., vazquez_martinez_evaluating_2023) risk underestimating the grammatical competence of models by failing to control for the influence of M{M}. In other words, a model could seemingly not differentiate between grammatical and ungrammatical strings, if the messages across grammatical and ungrammatical strings are not well-controlled. As we have argued here, this does not necessarily imply that the model has not learned generalizations about grammatical rules. If we are interested in isolating the model’s sensitivity to grammaticality, we have to use carefully designed procedures to factor out M{M}.
@@
One such procedure is minimal pair string comparisons, where the members of the minimal pair are closely matched in M{M} but are hypothesized to differ in G{G}. While controlled minimal pair comparisons are hardly new in NLP (e.g., marvin_targeted_2018; futrell_neural_2019), we have provided new theoretical grounding for these practices. In addition, our work lays the foundation for using computational techniques to isolate the effects of M{M} and G{G}, which has been explored in recent work (stanczak2023grammatical).
@@
Our analyses also raise new questions regarding LMs’ grammatical knowledge. The poor separation achieved by state-of-the-art LMs in Section˜6 feels counterintuitive: if LMs virtually always produce grammatical strings (under standard sampling procedures), then why is there so much overlap between the probabilities assigned to grammatical and ungrammatical strings? This tension between discriminative failures and generative abilities could be seen as a specific realization of the “generative AI paradox” (west_generative_2024), and also connects to recent work demonstrating that language identification is impossible except in highly constrained cases (gold_language_1967; angluin_inductive_1980), whereas language generation is possible for any countable list of languages (kleinberg_language_2024).
@@
leivada_reply_2024 argue that comparing isolated acceptability judgments in humans against minimal-pair probability differences in models is not comparing “apples with apples”, and thus unfair. We hope that our theoretical model can bring greater clarity to the issue of what makes a fair comparison. When comparing the (cognitive) abilities of two groups—humans and models, younger and older children, or even two different animal species—we maintain that the researcher must design assessments with that group’s (cognitive) computational architecture in mind. Applying the same evaluation method, which might impose auxiliary challenges on one group but not the other, can artificially increase apparent intergroup differences (firestone_performance_2020; lampinen_can_2024; hu_auxiliary_2024). For example: the intelligence of a squirrel should not be judged based on its ability to solve a Rubik’s cube. [^8] This example is due to Andrew Lampinen. Similarly, using metalinguistic judgments or isolated string probability as a window into grammaticality ignores the reality of what LMs are, and what they are trained to do—namely, to maximize the probability of strings from a corpus. More broadly, linguistic theory continues to suggest new ways to evaluate LMs, just as modern LMs provide new tools for studying the relationship between probability and grammaticality.